# Diffusion Models for Image Restoration and Enhancement: A Comprehensive Survey

### v2

@@ 4.1 Distortion Simulation
Notably, the real-world distortions are usually blind/u-nknown, where the distributions are different from the simple synthetic distortions. For supervised-learning-based IR, this requires the restoration network to have strong generalization capability or the synthetic datasets can cover the real-world distortions. From the causality perspective (Li et al., 2023b), these two purposes all rely on simulating diverse distortions that are similar to real-world distortions, which we called distortion simulation/augmentation. There are several representative DM-based IR methods (Sahak et al., 2023; Yin et al., 2022) utilizing the distortion simulations to improve the robustness of their methods for real-world degradations. The representative method is SR3+ (Sahak et al., 2023), which is based on the diffusion model from SR3 (Saharia et al., 2022b) and introduces the second-order degradations simulation of RealESRGAN (Wang et al., 2021b) for the training. Similarly, to simulate the real-world degradation, Yang et al.(Yang et al., 2023b) propose to synthesize the real-world distorted/clean training pairs using diffusion models, where the distorted images are initialized with second-order degradation in RealESRGAN (Wang et al., 2021b).
@@ 4.2 Kernel Estimation
Kernel Estimation is first proposed in blind image restoration (Tzikas et al., 2009; Huang and Xia, 2020; Vasu et al., 2018; Oliveira et al., 2013), where the degradation can be modeled as $y=(x*k)\downarrow_{s}+n$. Here, $k$ is the degradation kernel, and $n$ is the additive noise. Under this setting, the kernel $k$ can be estimated as the guidance to boost the adaptability of the restoration network. Inspired by this, BlindDPS (Chung et al., 2023a) and GibbsDDRM (Murata et al., 2023) attempt to solve the blind inverse problem by estimating the unknown degradation kernel in the sampling process. In particular, BlindDPS (Chung et al., 2023a) exploit the DPS (Chung et al., 2023b) architecture and exploit one parallel diffusion model for the degradation kernel estimation. The diffusion model for kernel estimation is pre-trained on synthetic kernels. Unlike BlindDPS, GibbsDDRM (Murata et al., 2023) achieves the sampling process with partially collapsed Gibbs sampler (Van Dyk and Park, 2008), which samples both kernel parameter and image together from the joint posterior $p(x_{t}|k,y)$.
@@ 4.3 Domain Translation
In the real world, it is hard to collect distorted/clean image pairs. Although some works attempt to simulate the degradation process of real distorted images, the distribution of the synthesized distortion is still far away from the real-world one. To further solve the Real-world IR problem, a series of works explore the domain translation techniques for image restoration. Domain Translation (Anoosheh et al., 2018; Lin et al., 2019; Chu and Wang, 2018; Saunders, 2022; Lugmayr et al., 2019b) aims to translate the image from one domain to another domain. From the domain translation perspective, synthetic distorted images, real-world distorted images, and high-quality images can be regarded as three different domains, that share the same contents.
The domain translation-based works for DM-based IR can be roughly divided into two categories: 1) The first line (Yang et al., 2023b) aims to simulate more reliable real-world distorted/clean image pairs by translating the low-quality images from the synthetic domain to the real-world domain. In this way, the simulated datasets can enable the restoration network with better restoration capability for real-world degradation. For instance, Yang et al.(Yang et al., 2023b) is the first to exploit the pre-trained diffusion model to synthesize real-world training pairs, where the diffusion model is pre-trained with real-world low-quality images, and the translation is achieved by warping the synthetic low-quality image to noise space (i.e., the generation inversion). 2) Another popular line (Wei et al., 2023) exploits unsupervised learning, where domain translation is achieved by the cycle consistency constraint. In particular, two generators construct one cycle path, where one generator aims to translate the distorted image to the distortion-free image, and another generator is utilized to translate the clean image to the distorted image. This enables unsupervised training with unpaired real-world distorted and high-quality images. As the representative work, RainDiffusion (Wei et al., 2023) proposes to remove the rain with two cooperative branches, where the non-diffusive translation branch aims to utilize the pre-trained cycle-consistent generators to produce initial paired clean/rainy images, and diffusive translation branch leverages the multi-scale diffusion models to refine the results.
@@ 4.4 Distortion-invariant Diffusion Model
Since the blind distortions are usually diverse and complicated, the diffusion model is required to own the generation capability for these distortions (i.e., the distortion-invariant capability) in the real world. To achieve the distortion-invariant diffusion model, DifFace et al.(Yue and Loy, 2024) introduces a pre-trained restoration network, e.g., SRCNN (Dong et al., 2014) or SwinIR (Liu et al., 2021), to obtain an initial clean image as a sampling start point $x_{N}$, where the restoration network is trained with the second-order degradation from RealE-SRGAN (Wang et al., 2021b), thereby exhibiting good generalization capability and produce the distortion invariant initial clean image for diffusion model. Ren et al.(Ren et al., 2022) proposes to achieve a distortion-invariant diffusion model with multi-scale degradation-invariant guidance information. They employ distortion augmentation strategies on degraded images to obtain invariant representation with structure information as guidance. In contrast, Wang (Wang et al., 2023i) exploits the low-pass filter to filter the distortion invariant components in the low-quality image since the different real-world distorted images usually share the same structure information. As shown in Fig. 11, they adopt a simple iterative refinement similar to ILVR during the sampling stage. After obtaining the degradation-invariant $\hat{x}_{0}$, they use an enhancement module (Powerful CNN-based or transformer-based restoration methods) to further improve the image quality.
Figure 11: The architecture of DR2 (Wang et al., 2023i). It consists degradation removal module and an enhancement module. The input image is first sent to a diffusion process to remove degradation information gradually. The output degradation-free image will be enhanced by a pre-trained face restoration network.
@@
Low-light enhancement. Real-world low-light image restoration aims to enhance and restore images captured under poor lighting conditions, which is susceptible to low brightness, high noise, color distortion, and reduced contrast, posing significant challenges for restoration methods. To achieve this, LLDiffusion (Wang et al., 2023e) introduces a degradation generation model to capture the degradation representation for low-light enhancement. This degradation-aware condition enhances the performance of diffusion-based low-light image restoration by providing more informative priors. To improve the efficiency of the diffusion model while preserving the generation capability, DiffLL (Jiang et al., 2023a) decouples the low-light image with wavelet-transform, using the diffusion model only to the low-frequency component. This approach significantly reduces the inference time while preserving the high-frequency details through a dedicated high-frequency restoration module(HFRM). PyDiffusion (Zhou et al., 2023) incorporates a multi-scale generation framework, applying the diffusion model at different scales to mitigate the computational inefficiency of direct high-resolution generation. This hierarchical approach enhances both efficiency and performance. There are also two diffusion-based works incorporating physically inspired priors to serve as conditioning inputs to the diffusion model, improving restoration quality by explicitly modeling the underlying lighting and reflectance properties. For instance, Diff-Retinex (Yi et al., 2023) decomposes low-light images into illumination and reflectance maps based on the Retinex model. LightenDiffusion (Jiang et al., 2024a) introduces an unsupervised training mechanism, it decouples the unpaired normal-light image and low-light images to extract the illumination prior from normal-light image and content reflection from low-light image, then combine them with diffusion model to refine it.
As shown in Table 1, we have compared recent diffusion-based works on real-world low-light image restoration on commonly-used synthetic dataset LOL (Wei et al., 2018) and three commonly-used real-world datasets, including DICM (Lee et al., 2013), NPE (Wang et al., 2013), and VV (Vonikakis et al., 2018). We can observe that DiffLL(Jiang et al., 2023a) and LightenDiffusion(Jiang et al., 2024a) achieve optimal generalization in terms of subjective metrics, i.e., NIQE and PI. However, due to its unsupervised training mechanism, LightenDiffusion exhibits lower objective quality compared to DiffLL (Jiang et al., 2023a).
@@
Models LOL (Wei et al., 2018) DICM (Lee et al., 2013) NPE (Wang et al., 2013) VV (Vonikakis et al., 2018)
@@
CLE Diffusion (Yin et al., 2023) 25.51 0.890 0.160 4.505 3.361 5.249 3.512 3.240 3.470
GDP (Fei et al., 2023) 15.90 0.542 0.337 4.358 3.552 4.032 3.097 4.683 3.431
DiffLL (Jiang et al., 2023a) 26.34 0.845 0.217 3.636 2.936 3.716 2.629 2.351 2.869
LLDiffusion (Wang et al., 2023e) 24.65 0.843 0.075 - - - - - -
Diff-Retinex (Yi et al., 2023) 21.98 0.863 0.048 4.361 3.394 4.996 3.392 3.087 3.350
PyDiffusion (Zhou et al., 2023) 23.28 0.859 0.108 4.499 3.792 4.082 3.268 4.360 3.678
LightenDiffusion (Jiang et al., 2024a) 20.45 0.803 0.192 3.724 3.144 3.618 2.879 2.941 2.558
@@
Adverse weather removal. As the pioneering work, WeathDiff (Özdenizci and Legenstein, 2023) shares the same diffusion architecture with SR3, but it can support arbitrary image restoration by introducing the mean estimated noise of overlapped patches in the reverse process. In contrast, T3-Diffusion (Chen et al., 2024b) defined a group of degradation-specific sub-prompts and utilizes the dynamical composition of them based on unseen degradation for weather removing, which outperforms DA-CLIP (Luo et al., 2024) and Weather-Diff. ReviveDiff (Huang et al., 2024) enhances the perception levels for various weather degradation by introducing the coarse and fine branches into the diffusion model. RainDiff/RainDiffusion (Shen et al., 2023) proposes an unpaired cycle consistency framework for image deraining, eliminating the need for paired training data, making it highly suitable for real-world applications. Depart from that, there are also two diffusion-based works for image dehazing. Dehaze-DDPM (Yu et al., 2023a) employs a two-stage dehazing pipeline, combining the Atmospheric Scattering Model (ASM) for physical modeling in the first stage and diffusion-based generative recovery of lost structural details in the second. Additionally, Wang et al. (Wang et al., 2023b) enhance dehazing by emphasizing mid-to-high-frequency components in the diffusion process, leveraging a frequency-spectrum filter and skip connections to inject high-frequency details into the U-Net architecture.
@@
Shadow removal. ShadowDiffusion (Guo et al., 2023a) leverages a pre-trained transformer backbone to extract the degradation prior (i.e., the degradation-related features) from the distorted reference image. This extracted degradation prior is exploited as the auxiliary to refine the generated shadow mask and serves as the condition for shadow-free image generation. DeS3 (Jin et al., 2024) introduces a novel adaptive attention mechanism to distinguish the underlying objects and shadow regions, which is mask-free compared with existing works on shadow removal. Moreover, it utilizes a pre-trained ViT backbone as loss to enhance the scene understanding capability. As described in DeS3, it can obtain a better performance compared with ShadowDiffusion.
@@
Potential limitations. The potential limitations of current diffusion-based methods in addressing real-world IR are as: (i) Limited task relevance. Some diffusion-based image restoration (IR) methods overlook the inherent characteristics of the tasks themselves and instead rely solely on dataset-specific fitting; (ii) Limited comparison reasonability. As discussed earlier, diffusion-based methods (Özdenizci and Legenstein, 2023; Wei et al., 2023) often use different datasets in some IR task, making it difficult to identify which approach is truly more effective. Furthermore, real-world datasets are challenging to collect and are often insufficient for comprehensive evaluation; (iii) Limited restoration efficiency. As stated in Section 6 of our manuscript, the impressive perceptual quality of diffusion-based IR methods is largely dependent on multiple sampling steps. Although one-step diffusion-based IR methods (Wang et al., 2024c; Wu et al., 2024a) have been proposed, accelerating the sampling process while maintaining restoration quality remains a significant challenge; (iv) Limited pixel-wise consistency. Diffusion-based real image restoration methods (Yu et al., 2024; Jiang et al., 2024a) share a common weakness with GAN-based approaches, i.e., they are susceptible to hallucinated textures due to the generative priors stored within the diffusion model. Addressing this issue remains a critical challenge.
@@ 5.1 Datasets and Implementation Details
RealESRGAN+ (Wang et al., 2021b) 28.18 0.776 0.115 13.57 24.37 0.734 0.123 20.29 0.0327 16.7M 73.43G 1
CAL-GAN (Park et al., 2023) 28.96 0.790 0.108 13.22 25.33 0.762 0.117 18.47 0.0327 16.7M 73.43G 1
SeD (Li et al., 2024a) 29.27 0.803 0.094 12.51 25.93 0.780 0.107 16.39 0.0327 16.7M 73.43G 1
RCoT (Tang et al., 2024) 29.63 0.819 0.279 20.04 24.87 0.742 0.270 33.63 0.0164 6.74M 204.74G 1
PromptIR (Potlapalli et al., 2023) 30.73 0.844 0.262 16.75 26.71 0.804 0.209 26.41 0.0602 33.19M 11.91G 1
Restormer (Zamir et al., 2022) 30.73 0.844 0.261 16.65 26.74 0.804 0.209 26.11 0.0557 26.32M 10.84G 1
RCAN (Zhang et al., 2018b) 30.78 0.844 0.255 16.41 26.74 0.807 0.196 24.23 0.0863 15.4M 65.25G 1
IPT (Chen et al., 2021) 30.92 0.847 0.253 15.76 27.33 0.820 0.189 22.92 0.0379 115M 43.41G 1
SwinIR (Liu et al., 2021) 31.08 0.852 0.247 15.29 27.46 0.826 0.184 21.98 0.0384 11.8M 37.17G 1
HAT (Chen et al., 2023a) 31.22 0.855 0.245 15.33 27.97 0.836 0.173 20.49 0.0613 20.6M 102.4G 1
MambaIRv2 (Guo et al., 2024a) 31.24 0.858 0.246 15.94 28.07 0.838 0.171 20.83 0.0626 34.2M 82.3G 1
SR3 (Saharia et al., 2022b) 20.37 0.629 0.370 46.04 20.60 0.635 0.288 43.55 50.4 155.3M 155.2G 100
SRDiff (Li et al., 2022b) 28.62 0.789 0.130 16.93 25.14 0.758 0.138 20.65 4.5 13.2M 84.22G 100
Resshift (Yue et al., 2023) 29.38 0.805 0.130 15.41 25.68 0.766 0.122 20.19 - - - 15
DiffBIR (Lin et al., 2024) 25.65 0.702 0.187 26.45 23.49 0.676 0.204 37.68 4.65 1716.71M 189.47G 50
SUPIR (Yu et al., 2024) 27.19 0.728 0.192 24.94 23.26 0.669 0.182 33.38 7.64 4801.1M 214.43G 50
@@
RealESRGAN (Wang et al., 2021b) 25.94 0.705 0.123 30.29 29.48 0.801 0.099 26.60 0.0327 73.43G 1
CAL-GAN (Park et al., 2023) 26.92 0.736 0.122 29.14 30.00 0.815 0.101 22.40 0.0327 73.43G 1
SeD (Li et al., 2024a) 27.01 0.741 0.109 27.10 29.62 0.807 0.089 18.96 0.0327 73.43G 1
RCOT (Tang et al., 2024) 27.36 0.755 0.297 57.42 30.28 0.825 0.229 75.13 0.0164 204.74G 1
PromptIR (Potlapalli et al., 2023) 28.35 0.787 0.274 50.13 30.99 0.842 0.224 62.79 0.0602 11.91G 1
Restormer (Zamir et al., 2022) 28.34 0.786 0.274 49.78 30.98 0.842 0.224 62.23 0.0557 10.84G 1
RCAN (Zhang et al., 2018b) 28.35 0.786 0.269 50.92 30.95 0.839 0.222 60.87 0.0863 65.25G 1
IPT (Chen et al., 2021) 28.39 0.787 0.268 50.26 30.98 0.839 0.220 59.47 0.0379 43.41G 1
SwinIR (Liu et al., 2021) 28.56 0.792 0.263 49.57 31.03 0.841 0.220 60.72 0.0384 37.17G 1
HAT (Chen et al., 2023a) 28.67 0.795 0.262 48.20 31.08 0.841 0.220 60.69 0.0613 102.4G 1
MambaIR (Guo et al., 2024b) 28.62 0.793 0.264 48.87 31.07 0.843 0.219 59.36 0.0626 82.30G 1
ILVR (Choi et al., 2021) 27.40 0.871 0.21 43.76 31.59 0.878 0.22 32.24 41.3 1113.75 250
SNIPS (Kawar et al., 2021) 24.31 0.684 0.21 124.54 27.34 0.675 0.27 105.19 31.4 - 1000
DDRM (Kawar et al., 2022a) 27.38 0.869 0.22 40.75 31.64 0.946 0.19 31.04 10.1 1113.75 20
DPS (Chung et al., 2023b) 25.88 0.814 0.15 39.24 26.95 0.878 0.18 29.97 141.2 1113.75 1000
DDNM (Wang et al., 2023g) 27.46 0.871 0.15 40.15 31.64 0.945 0.16 28.27 15.5 1113.75 100
GDP (Fei et al., 2023) 26.51 0.832 0.14 38.45 28.65 0.876 0.17 27.51 3.1 1113.76 20

### v3

@@ 4.1 Distortion Simulation
Notably, the real-world distortions are usually blind/u-nknown, where the distributions are different from the simple synthetic distortions. For supervised-learning-based IR, this requires the restoration network to have strong generalization capability or the synthetic datasets can cover the real-world distortions. From the causality perspective (li2023learningcausalIR, ), these two purposes all rely on simulating diverse distortions that are similar to real-world distortions, which we called distortion simulation/augmentation. There are several representative DM-based IR methods (SR3+, ; yin2022diffgar, ) utilizing the distortion simulations to improve the robustness of their methods for real-world degradations. The representative method is SR3+ (SR3+, ), which is based on the diffusion model from SR3 (SR3, ) and introduces the second-order degradations simulation of RealESRGAN (wang2021realesrgan, ) for the training. Similarly, to simulate the real-world degradation, Yang et al.(yang2023synthesizing, ) propose to synthesize the real-world distorted/clean training pairs using diffusion models, where the distorted images are initialized with second-order degradation in RealESRGAN (wang2021realesrgan, ).
@@ 4.2 Kernel Estimation
Kernel Estimation is first proposed in blind image restoration (tzikas2009variational-blind0, ; huang2020joint-blind1, ; vasu2018non-blind2, ; oliveira2013parametric-blind3, ), where the degradation can be modeled as $y=(x*k)\downarrow_{s}+n$. Here, $k$ is the degradation kernel, and $n$ is the additive noise. Under this setting, the kernel $k$ can be estimated as the guidance to boost the adaptability of the restoration network. Inspired by this, BlindDPS (blinddps, ) and GibbsDDRM (murata2023gibbsddrm, ) attempt to solve the blind inverse problem by estimating the unknown degradation kernel in the sampling process. In particular, BlindDPS (blinddps, ) exploit the DPS (dps, ) architecture and exploit one parallel diffusion model for the degradation kernel estimation. The diffusion model for kernel estimation is pre-trained on synthetic kernels. Unlike BlindDPS, GibbsDDRM (murata2023gibbsddrm, ) achieves the sampling process with partially collapsed Gibbs sampler (van2008partially-sampler, ), which samples both kernel parameter and image together from the joint posterior $p(x_{t}|k,y)$.
@@ 4.3 Domain Translation
In the real world, it is hard to collect distorted/clean image pairs. Although some works attempt to simulate the degradation process of real distorted images, the distribution of the synthesized distortion is still far away from the real-world one. To further solve the Real-world IR problem, a series of works explore the domain translation techniques for image restoration. Domain Translation (anoosheh2018combogan-translation0, ; lin2019learning-translation0.5, ; chu2018survey-translation1, ; saunders2022domain-translation2, ; lugmayr2019unsupervised-translation, ) aims to translate the image from one domain to another domain. From the domain translation perspective, synthetic distorted images, real-world distorted images, and high-quality images can be regarded as three different domains, that share the same contents.
The domain translation-based works for DM-based IR can be roughly divided into two categories: 1) The first line (yang2023synthesizing, ) aims to simulate more reliable real-world distorted/clean image pairs by translating the low-quality images from the synthetic domain to the real-world domain. In this way, the simulated datasets can enable the restoration network with better restoration capability for real-world degradation. For instance, Yang et al.(yang2023synthesizing, ) is the first to exploit the pre-trained diffusion model to synthesize real-world training pairs, where the diffusion model is pre-trained with real-world low-quality images, and the translation is achieved by warping the synthetic low-quality image to noise space (i.e., the generation inversion). 2) Another popular line (wei2023raindiffusion, ) exploits unsupervised learning, where domain translation is achieved by the cycle consistency constraint. In particular, two generators construct one cycle path, where one generator aims to translate the distorted image to the distortion-free image, and another generator is utilized to translate the clean image to the distorted image. This enables unsupervised training with unpaired real-world distorted and high-quality images. As the representative work, RainDiffusion (wei2023raindiffusion, ) proposes to remove the rain with two cooperative branches, where the non-diffusive translation branch aims to utilize the pre-trained cycle-consistent generators to produce initial paired clean/rainy images, and diffusive translation branch leverages the multi-scale diffusion models to refine the results.
@@ 4.4 Distortion-invariant Diffusion Model
Since the blind distortions are usually diverse and complicated, the diffusion model is required to own the generation capability for these distortions (i.e., the distortion-invariant capability) in the real world. To achieve the distortion-invariant diffusion model, DifFace et al.(yue2022difface, ) introduces a pre-trained restoration network, e.g., SRCNN (SR1_SRCNN-cnn, ) or SwinIR (swinIR-transformer, ), to obtain an initial clean image as a sampling start point $x_{N}$, where the restoration network is trained with the second-order degradation from RealE-SRGAN (wang2021realesrgan, ), thereby exhibiting good generalization capability and produce the distortion invariant initial clean image for diffusion model. Ren et al.(dg-dpm, ) proposes to achieve a distortion-invariant diffusion model with multi-scale degradation-invariant guidance information. They employ distortion augmentation strategies on degraded images to obtain invariant representation with structure information as guidance. In contrast, Wang (wang2023dr2, ) exploits the low-pass filter to filter the distortion invariant components in the low-quality image since the different real-world distorted images usually share the same structure information. As shown in Fig. 11, they adopt a simple iterative refinement similar to ILVR during the sampling stage. After obtaining the degradation-invariant $\hat{x}_{0}$, they use an enhancement module (Powerful CNN-based or transformer-based restoration methods) to further improve the image quality.
Figure 11: The architecture of DR2 (wang2023dr2, ). It consists degradation removal module and an enhancement module. The input image is first sent to a diffusion process to remove degradation information gradually. The output degradation-free image will be enhanced by a pre-trained face restoration network.
@@
Low-light enhancement. Real-world low-light image restoration aims to enhance and restore images captured under poor lighting conditions, which is susceptible to low brightness, high noise, color distortion, and reduced contrast, posing significant challenges for restoration methods. To achieve this, LLDiffusion (wang2023lldiffusion-LLDiffuion, ) introduces a degradation generation model to capture the degradation representation for low-light enhancement. This degradation-aware condition enhances the performance of diffusion-based low-light image restoration by providing more informative priors. To improve the efficiency of the diffusion model while preserving the generation capability, DiffLL (jiang2023lowDiffLL, ) decouples the low-light image with wavelet-transform, using the diffusion model only to the low-frequency component. This approach significantly reduces the inference time while preserving the high-frequency details through a dedicated high-frequency restoration module(HFRM). PyDiffusion (zhou2023pyramid-pydiff, ) incorporates a multi-scale generation framework, applying the diffusion model at different scales to mitigate the computational inefficiency of direct high-resolution generation. This hierarchical approach enhances both efficiency and performance. There are also two diffusion-based works incorporating physically inspired priors to serve as conditioning inputs to the diffusion model, improving restoration quality by explicitly modeling the underlying lighting and reflectance properties. For instance, Diff-Retinex (yi2023diff-retinex, ) decomposes low-light images into illumination and reflectance maps based on the Retinex model. LightenDiffusion (jiang2024lightendiffusion, ) introduces an unsupervised training mechanism, it decouples the unpaired normal-light image and low-light images to extract the illumination prior from normal-light image and content reflection from low-light image, then combine them with diffusion model to refine it.
As shown in Table 1, we have compared recent diffusion-based works on real-world low-light image restoration on commonly-used synthetic dataset LOL (lol, ) and three commonly-used real-world datasets, including DICM (lee2013contrastDICM, ), NPE (wang2013naturalnessNPE, ), and VV (vonikakis2018evaluationVV, ). We can observe that DiffLL(jiang2023lowDiffLL, ) and LightenDiffusion(jiang2024lightendiffusion, ) achieve optimal generalization in terms of subjective metrics, i.e., NIQE and PI. However, due to its unsupervised training mechanism, LightenDiffusion exhibits lower objective quality compared to DiffLL (jiang2023lowDiffLL, ).
@@
Models LOL (lol, ) DICM (lee2013contrastDICM, ) NPE (wang2013naturalnessNPE, ) VV (vonikakis2018evaluationVV, )
@@
CLE Diffusion (yin2023cle-CLE-Diffusion, ) 25.51 0.890 0.160 4.505 3.361 5.249 3.512 3.240 3.470
GDP (GDP, ) 15.90 0.542 0.337 4.358 3.552 4.032 3.097 4.683 3.431
DiffLL (jiang2023lowDiffLL, ) 26.34 0.845 0.217 3.636 2.936 3.716 2.629 2.351 2.869
LLDiffusion (wang2023lldiffusion-LLDiffuion, ) 24.65 0.843 0.075 - - - - - -
Diff-Retinex (yi2023diff-retinex, ) 21.98 0.863 0.048 4.361 3.394 4.996 3.392 3.087 3.350
PyDiffusion (zhou2023pyramid-pydiff, ) 23.28 0.859 0.108 4.499 3.792 4.082 3.268 4.360 3.678
LightenDiffusion (jiang2024lightendiffusion, ) 20.45 0.803 0.192 3.724 3.144 3.618 2.879 2.941 2.558
@@
Adverse weather removal. As the pioneering work, WeathDiff (weather-diff, ) shares the same diffusion architecture with SR3, but it can support arbitrary image restoration by introducing the mean estimated noise of overlapped patches in the reverse process. In contrast, T3-Diffusion (chen2024teachingt3diffusion, ) defined a group of degradation-specific sub-prompts and utilizes the dynamical composition of them based on unseen degradation for weather removing, which outperforms DA-CLIP (luo2023controlling-DACLIP, ) and Weather-Diff. ReviveDiff (huang2024revivediff, ) enhances the perception levels for various weather degradation by introducing the coarse and fine branches into the diffusion model. RainDiff/RainDiffusion (shen2023rethinkingRainDiff, ) proposes an unpaired cycle consistency framework for image deraining, eliminating the need for paired training data, making it highly suitable for real-world applications. Depart from that, there are also two diffusion-based works for image dehazing. Dehaze-DDPM (yu2023high-HazeDDPM, ) employs a two-stage dehazing pipeline, combining the Atmospheric Scattering Model (ASM) for physical modeling in the first stage and diffusion-based generative recovery of lost structural details in the second. Additionally, Wang et al. (wang2023frequency-dehazy, ) enhance dehazing by emphasizing mid-to-high-frequency components in the diffusion process, leveraging a frequency-spectrum filter and skip connections to inject high-frequency details into the U-Net architecture.
@@
Shadow removal. ShadowDiffusion (guo2023shadowdiffusion, ) leverages a pre-trained transformer backbone to extract the degradation prior (i.e., the degradation-related features) from the distorted reference image. This extracted degradation prior is exploited as the auxiliary to refine the generated shadow mask and serves as the condition for shadow-free image generation. DeS3 (jin2022des3, ) introduces a novel adaptive attention mechanism to distinguish the underlying objects and shadow regions, which is mask-free compared with existing works on shadow removal. Moreover, it utilizes a pre-trained ViT backbone as loss to enhance the scene understanding capability. As described in DeS3, it can obtain a better performance compared with ShadowDiffusion.
@@
Potential limitations. The potential limitations of current diffusion-based methods in addressing real-world IR are as: (i) Limited task relevance. Some diffusion-based image restoration (IR) methods overlook the inherent characteristics of the tasks themselves and instead rely solely on dataset-specific fitting; (ii) Limited comparison reasonability. As discussed earlier, diffusion-based methods (weather-diff, ; wei2023raindiffusion, ) often use different datasets in some IR task, making it difficult to identify which approach is truly more effective. Furthermore, real-world datasets are challenging to collect and are often insufficient for comprehensive evaluation; (iii) Limited restoration efficiency. As stated in Section 6 of our manuscript, the impressive perceptual quality of diffusion-based IR methods is largely dependent on multiple sampling steps. Although one-step diffusion-based IR methods (wang2023sinsr, ; wu2024oneOSEDiff, ) have been proposed, accelerating the sampling process while maintaining restoration quality remains a significant challenge; (iv) Limited pixel-wise consistency. Diffusion-based real image restoration methods (yu2024scaling-SUPIR, ; jiang2024lightendiffusion, ) share a common weakness with GAN-based approaches, i.e., they are susceptible to hallucinated textures due to the generative priors stored within the diffusion model. Addressing this issue remains a critical challenge.
@@ 5.1 Datasets and Implementation Details
RealESRGAN+ (wang2021realesrgan, ) 28.18 0.776 0.115 13.57 24.37 0.734 0.123 20.29 0.0327 16.7M 73.43G 1
CAL-GAN (CAL-GAN, ) 28.96 0.790 0.108 13.22 25.33 0.762 0.117 18.47 0.0327 16.7M 73.43G 1
SeD (li2024sed, ) 29.27 0.803 0.094 12.51 25.93 0.780 0.107 16.39 0.0327 16.7M 73.43G 1
RCoT (tang2024residualRCoT, ) 29.63 0.819 0.279 20.04 24.87 0.742 0.270 33.63 0.0164 6.74M 204.74G 1
PromptIR (potlapalli2023promptir, ) 30.73 0.844 0.262 16.75 26.71 0.804 0.209 26.41 0.0602 33.19M 11.91G 1
Restormer (zamir2022restormer_deblur-transformer, ) 30.73 0.844 0.261 16.65 26.74 0.804 0.209 26.11 0.0557 26.32M 10.84G 1
RCAN (zhang2018image-RCAN, ) 30.78 0.844 0.255 16.41 26.74 0.807 0.196 24.23 0.0863 15.4M 65.25G 1
IPT (chen2021preIPT, ) 30.92 0.847 0.253 15.76 27.33 0.820 0.189 22.92 0.0379 115M 43.41G 1
SwinIR (swinIR-transformer, ) 31.08 0.852 0.247 15.29 27.46 0.826 0.184 21.98 0.0384 11.8M 37.17G 1
HAT (chen2023hat, ) 31.22 0.855 0.245 15.33 27.97 0.836 0.173 20.49 0.0613 20.6M 102.4G 1
MambaIRv2 (guo2024mambairv2, ) 31.24 0.858 0.246 15.94 28.07 0.838 0.171 20.83 0.0626 34.2M 82.3G 1
SR3 (SR3, ) 20.37 0.629 0.370 46.04 20.60 0.635 0.288 43.55 50.4 155.3M 155.2G 100
SRDiff (li2022srdiff, ) 28.62 0.789 0.130 16.93 25.14 0.758 0.138 20.65 4.5 13.2M 84.22G 100
Resshift (yue2023resshift, ) 29.38 0.805 0.130 15.41 25.68 0.766 0.122 20.19 - - - 15
DiffBIR (lin2023diffbir, ) 25.65 0.702 0.187 26.45 23.49 0.676 0.204 37.68 4.65 1716.71M 189.47G 50
SUPIR (yu2024scaling-SUPIR, ) 27.19 0.728 0.192 24.94 23.26 0.669 0.182 33.38 7.64 4801.1M 214.43G 50
@@
RealESRGAN (wang2021realesrgan, ) 25.94 0.705 0.123 30.29 29.48 0.801 0.099 26.60 0.0327 73.43G 1
CAL-GAN (CAL-GAN, ) 26.92 0.736 0.122 29.14 30.00 0.815 0.101 22.40 0.0327 73.43G 1
SeD (li2024sed, ) 27.01 0.741 0.109 27.10 29.62 0.807 0.089 18.96 0.0327 73.43G 1
RCOT (tang2024residualRCoT, ) 27.36 0.755 0.297 57.42 30.28 0.825 0.229 75.13 0.0164 204.74G 1
PromptIR (potlapalli2023promptir, ) 28.35 0.787 0.274 50.13 30.99 0.842 0.224 62.79 0.0602 11.91G 1
Restormer (zamir2022restormer_deblur-transformer, ) 28.34 0.786 0.274 49.78 30.98 0.842 0.224 62.23 0.0557 10.84G 1
RCAN (zhang2018image-RCAN, ) 28.35 0.786 0.269 50.92 30.95 0.839 0.222 60.87 0.0863 65.25G 1
IPT (chen2021preIPT, ) 28.39 0.787 0.268 50.26 30.98 0.839 0.220 59.47 0.0379 43.41G 1
SwinIR (swinIR-transformer, ) 28.56 0.792 0.263 49.57 31.03 0.841 0.220 60.72 0.0384 37.17G 1
HAT (chen2023hat, ) 28.67 0.795 0.262 48.20 31.08 0.841 0.220 60.69 0.0613 102.4G 1
MambaIR (guo2024mambair, ) 28.62 0.793 0.264 48.87 31.07 0.843 0.219 59.36 0.0626 82.30G 1
ILVR (choi2021ilvr, ) 27.40 0.871 0.21 43.76 31.59 0.878 0.22 32.24 41.3 1113.75 250
SNIPS (kawar2021snips, ) 24.31 0.684 0.21 124.54 27.34 0.675 0.27 105.19 31.4 - 1000
DDRM (DDRMDRM, ) 27.38 0.869 0.22 40.75 31.64 0.946 0.19 31.04 10.1 1113.75 20
DPS (dps, ) 25.88 0.814 0.15 39.24 26.95 0.878 0.18 29.97 141.2 1113.75 1000
DDNM (ddnm, ) 27.46 0.871 0.15 40.15 31.64 0.945 0.16 28.27 15.5 1113.75 100
GDP (GDP, ) 26.51 0.832 0.14 38.45 28.65 0.876 0.17 27.51 3.1 1113.76 20